OpenStudy (anonymous):

simplify the expression (-6+i)/(-5+i)

OpenStudy (anonymous):

The options are \[31+i \div 26\] \[29+i \div 26\] \[31+i \div 24\] \[31+11i \div 26\]

OpenStudy (michele_laino):

please, you have to multiply both numerator and denominator by (-5-i), namely: \[\frac{( -6+i) }{ (-5+i) }=\frac{ (-6+i)*(-5-i) }{ (-5+i)*(-5-i) }=...\] please try to complete the above calculus

OpenStudy (anonymous):

I don't quite understand

OpenStudy (michele_laino):

@iatethepotato for example: \[(-6+i)*(-5-i)=-6*(-5-i)+i*(-5-i)=\] \[=30+6i-5i+1=31+i\] now, please do the same with the denominator

OpenStudy (anonymous):

the trick here is to get rid of the (i) in the denominator

OpenStudy (anonymous):

you have (-5+i) so multiply by (-5+i)/(-5+i)

OpenStudy (anonymous):

the denominator (-5+i)*(-5-i)=25+5i-5i-i^2=25-(-1)=25+1=26

OpenStudy (anonymous):

Okay I see now thank you :)

OpenStudy (anonymous):

YW